OpenStudy (anonymous):

Solve the following system. x - y = 10 x + y = 8 Answer (9, -1) (9, 1) (-1, 9) (-9, 1)

OpenStudy (bahrom7893):

Okay this is how u solve systems. First solve one equation for one variable.

OpenStudy (bahrom7893):

So solve the first equation for x, or y, doesn't matter. Let's solve for x: x-y=10 Add y to both sides. What do u get?

OpenStudy (anonymous):

x = y + 10

OpenStudy (bahrom7893):

Correct. Now plug that into the second equation: x + y = 8, but now that u know x, plug that instead of x in here. What do u get?

OpenStudy (anonymous):

2y+10=8

OpenStudy (bahrom7893):

correct. Subtract ten from both sides. What do u get?

OpenStudy (anonymous):

y = -1

OpenStudy (bahrom7893):

awesome. Now plug that back into the equation you had before.

OpenStudy (anonymous):

x + -1 = 8 x = 9 danks schun!

OpenStudy (bahrom7893):

Yup. Your points are given in the form (x; y). Find whichever's yours.

OpenStudy (bahrom7893):

good job lol.

OpenStudy (anonymous):

lol